OpenStudy (anonymous):

Calculate the degrees of freedom for a one sample t-test with 2822 US adults between the ages of 15 and 64 in the sample

OpenStudy (anonymous):

DF= n-1, n=2822

OpenStudy (anonymous):

That's all you have to do?

OpenStudy (anonymous):

Yeah, it's the total number of samples. I don't remember the full equation of the t-test, but that's what degree of freedom is

OpenStudy (anonymous):

So even if theres say 372 men in a sample and theres 128 million men in the population, the df would still be df= n-1 so df=372-1= 371

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes, because it's only one group. Have to correct myself, it's DF=n- (number of groups being tested) so, since it's between 15 and 64, that's 2. so it would be 2822-2 = 2820

OpenStudy (anonymous):

Ohh, you're a genius! I don't know how old you are but you just helped me so much out in my grad school level stats class

OpenStudy (anonymous):

I'm a junior bio major. We have to use stats a lot and just learned the t-test last summer.

OpenStudy (anonymous):

ahh, well stay at that school! they are way preparing you! i graduated with a communications degree and now going for Masters of Leadership and we're learning this supposedly to help out in studies.

OpenStudy (anonymous):

I'm thinking these people are crazy!!!

OpenStudy (anonymous):

Wow. Yeah, bio majors have to take a statistics class, but mine was YEARS ago and I don't remember anything. I learned it in one of my bio classes because of our experiments.

OpenStudy (anonymous):

Well, good luck! I'm hoping someone will help me with my question. It's calculus, lol.

OpenStudy (anonymous):

tell me about it, I did PSO (college in high school) and the Math they put me in was Math for Liberal Art majors... im 22 now and that was when I was like 18... so now you can understand why I was googling HELP WITH STATS HOMEWORK!!!!